# Open-Set Domain Adaptation Under Background Distribution Shift: Challenges and A Provably Efficient Solution

### v2

@@ A.5.2 Hyper-Parameters and Training
For all the methods, we keep the set of candidate target recall values consistent through all the experiments $\bm{\alpha}=[0.02,0.05,0.10,0.15,0.20,0.25,0.30,0.35,0.40,0.45]$. We consistently set the FPR threshold $\beta=0.01$ without optimizing it at all based on validation dataset. For each novel class identity, we repeat the experiments for 5 different randomly generated splits between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$ adhering to the definition and assumptions of background shift. For CIFAR100, we use ResNet18 backbone followed by a linear layer for classification and train the whole model from scratch for all the methods. For Amazon Reviews dataset, we use pretrained RoBERTa features followed by 2 linear layers for classification. For Amazon Reviews, due to computational limitations we resort to linear probing rather than finetuning the whole model and only finetune the last two linear layers. The primary hyperparameters we tune for stable convergence are learning rate, L2 weight penalty scalar to prevent overfitting, logit multiplier values that act as temperature controllers for softmax/sigmoid scores and gradient clipping to avoid exploding gradients. These hyper parameters are tuned based on a sample training and validation sets but are kept constant throughout the dataset and baseline across different seed values and novelty cases. Furthermore, the methods CoNoC & CoLOR require additional hyperparameters like dual learning rate and lagrange multipliers. Each output node is associated with Lagrange multipliers, which address a distinct primal-dual optimization problem owing to diverse target recall constraints. These multipliers are initialized to 1.0, while the dual learning rate is meticulously calibrated for CIFAR100 & Amazon Reviews datasets individually to ensure stable learning dynamics conducive to minimizing both the objective surrogate loss function for FPR and recall inequality constraints. Note that for ZOC, We used cosine similarity between image embeddings and the known class text embeddings to obtain the closed-set class predictions. Table 5 displays all the hyperparameters used for each of the baselines and datasets.
@@ A.6 Performance comparison based on average relative & absolute AU-ROC and AU-PRC scores
Table 6: SUN397 dataset with distribution shift due to varying proportions of subtypes of scenes/places. All the methods here use ResNet50 backbone pretrained on ImageNet1K_V1 Russakovsky2015imagenet. AUROC and AUPRC represent the performance for the novel category detection task while OSCR measures overall performance of the methods on both known and unknown classes. $\alpha$ is the mixture proportion column for the respective novel classes.
@@
Table 7: SUN397 dataset without distribution shift due to varying proportions of subtypes of scenes/places. All the methods here use ResNet50 backbone pretrained on ImageNet1K_V1 Russakovsky2015imagenet. AUROC and AUPRC represent the performance for the novel category detection task while OSCR measures overall performance of the methods on both known and unknown classes. $\alpha$ is the mixture proportion column for the respective novel classes.
@@
Table 8: SUN397 dataset with distribution shift due to varying proportions of subtypes of scenes/places. All the principled methods ($D$, uPU, nnPU, BODA & CoLOR) here use pretrained CLIP ViT-L/14 backbone from radford2021clip.
@@
Table 9: SUN397 dataset without any intended distribution shift. All the principled methods ($D$, uPU, nnPU, BODA & CoLOR) here use pretrained CLIP ViT-L/14 backbone from radford2021clip.
@@
Appendix B Analysis of toy example

### v3

@@ A.5.2 Hyper-Parameters and Training
We consistently set the FPR threshold $\beta=0.01$ without optimizing it at all based on validation dataset. For each novel class identity, we repeat the experiments for 5 different randomly generated splits between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$ adhering to the definition and assumptions of background shift. For CIFAR100, we use ResNet18 backbone followed by a linear layer for classification and train the whole model from scratch for all the methods. For Amazon Reviews dataset, we use pretrained RoBERTa features followed by 2 linear layers for classification. For Amazon Reviews, due to computational limitations we resort to linear probing rather than finetuning the whole model and only finetune the last two linear layers. The primary hyperparameters we tune for stable convergence are learning rate, L2 weight penalty scalar to prevent overfitting, logit multiplier values that act as temperature controllers for softmax/sigmoid scores and gradient clipping to avoid exploding gradients. These hyper parameters are tuned based on a sample training and validation sets but are kept constant throughout the dataset and baseline across different seed values and novelty cases. Furthermore, the methods CoNoC & CoLOR require additional hyperparameters like dual learning rate and lagrange multipliers. Each output node is associated with Lagrange multipliers, which address a distinct primal-dual optimization problem owing to diverse target recall constraints. These multipliers are initialized to 1.0, while the dual learning rate is meticulously calibrated for CIFAR100 & Amazon Reviews datasets individually to ensure stable learning dynamics conducive to minimizing both the objective surrogate loss function for FPR and recall inequality constraints. Note that for ZOC, We used cosine similarity between image embeddings and the known class text embeddings to obtain the closed-set class predictions. Table 5 displays all the hyperparameters used for each of the baselines and datasets.
A.5.3 Impact of search grid range and density on the performance
For all the methods, we keep the search grid of candidate target recall values consistent through all the experiments i.e. $\bm{\alpha}=[0.02,0.05,0.10,0.15,0.20,0.25,0.30,0.35,0.40,0.45]$.
Choice of the search grid range:
$\bm{\alpha}=[0.02,0.45]$
is not claimed to be universal, but chosen to study the most informative regime. We observed that when the size of the novel class (true size, not the one estimated by the model) is below $1\%$, all baselines failed. Accordingly, we set the lower bound of our search grid on that scale, but we show in experiments that including lower values does not change performance by much. As for the upper bound, we observe that our method is relatively robust when the true novel class size is small (i.e. the model wouldn’t choose a larger alpha if it had the possibility). We cap the upper bound of the gris at $0.45$ because when the novel class constitutes a large fraction of the target data, novelty detection becomes relatively easier. In such regimes, even simple inspection of random data points or sampling strategies can easily detect novel instances, enabling labeling and the use of alternative solutions. Our primary focus is therefore on the more challenging and practically relevant regime where novel classes are relatively rare (typically between 0.02 and 0.2). We further confirm this through our ablation study in table˜20. We note that this lower bound is not fundamental and may depend on factors such as dataset characteristics, feature representations, and model architecture. In practice, we recommend determining an appropriate $\bm{\alpha}$ range via simulation: known classes can be temporarily treated as unknown, and their detectability can be evaluated under varying assumed novel class ratios. This procedure provides empirical guidance for both the bottom value and density of the grid range.
Choice of search grid density: We added experiments on the SUN397 dataset using a ResNet50 backbone pretrained on ImageNet-1K (Table 19 and Figure 5 in the revised draft). These results show that CoLOR is performant for a fixed $\bm{\alpha}$ range with only minor performance variations across varying grid densities. This indicates that the method does not require finely tuned grids to achieve strong performance, and that the chosen density is sufficient to cover the relevant operating regime.
@@ A.6 Performance comparison based on average relative & absolute AU-ROC and AU-PRC scores
Table 6: SUN397 dataset with distribution shift due to varying proportions of subtypes of scenes/places. All the methods here use ResNet50 backbone pretrained on ImageNet1K_V1 Russakovsky et al. (2015). AUROC and AUPRC represent the performance for the novel category detection task while OSCR measures overall performance of the methods on both known and unknown classes. $\alpha$ is the mixture proportion column for the respective novel classes.
@@
Table 7: SUN397 dataset without distribution shift due to varying proportions of subtypes of scenes/places. All the methods here use ResNet50 backbone pretrained on ImageNet1K_V1 Russakovsky et al. (2015). AUROC and AUPRC represent the performance for the novel category detection task while OSCR measures overall performance of the methods on both known and unknown classes. $\alpha$ is the mixture proportion column for the respective novel classes.
@@
Table 8: SUN397 dataset with distribution shift due to varying proportions of subtypes of scenes/places. All the principled methods ($D$, uPU, nnPU, BODA & CoLOR) here use pretrained CLIP ViT-L/14 backbone from Radford et al. (2021).
@@
Table 9: SUN397 dataset without any intended distribution shift. All the principled methods ($D$, uPU, nnPU, BODA & CoLOR) here use pretrained CLIP ViT-L/14 backbone from Radford et al. (2021).
@@
[…]
Performance Metrics across 3 novel class groups.
$\bm{\alpha}$ AUROC AUPRC OSCR Top-1 Acc.
density 0 1 2 0 1 2 0 1 2 0 1 2
2 0.98 0.99 0.98 0.90 0.96 0.94 0.79 0.83 0.79 0.81 0.83 0.81
3 0.98 0.99 0.98 0.93 0.96 0.94 0.81 0.84 0.80 0.82 0.85 0.82
4 0.98 0.99 0.98 0.94 0.97 0.94 0.81 0.86 0.80 0.83 0.86 0.82
5 0.98 0.99 0.98 0.93 0.97 0.95 0.81 0.84 0.81 0.83 0.84 0.83
6 0.99 0.99 0.98 0.94 0.97 0.91 0.83 0.85 0.75 0.84 0.86 0.76
7 0.98 0.99 0.98 0.93 0.97 0.94 0.82 0.86 0.80 0.83 0.87 0.82
8 0.99 0.99 0.98 0.95 0.97 0.94 0.84 0.86 0.81 0.85 0.87 0.83
9 0.99 0.99 0.99 0.95 0.97 0.95 0.83 0.85 0.81 0.84 0.86 0.83
10 0.98 0.99 0.99 0.93 0.96 0.95 0.82 0.83 0.83 0.84 0.84 0.85
[…]
Performance Metrics across 3 novel class groups.
$\bm{\alpha}$ AUROC AUPRC OSCR Top-1 Acc.
range 0 1 2 0 1 2 0 1 2 0 1 2
$\left[0.001,0.01\right]$ 0.98 0.99 0.98 0.90 0.96 0.94 0.79 0.83 0.79 0.81 0.83 0.81
$\left[0.01,0.1\right]$ 0.98 0.99 0.98 0.93 0.96 0.94 0.81 0.84 0.80 0.82 0.85 0.82
$\left[0.1,1.0\right]$ 0.98 0.99 0.98 0.94 0.97 0.94 0.81 0.86 0.80 0.83 0.86 0.82
$\left[0.2,0.45\right]$ 0.97 0.98 0.99 0.59 0.87 0.95 0.72 0.77 0.83 0.75 0.79 0.85
$\left[0.3,0.45\right]$ 0.98 0.99 0.99 0.70 0.95 0.94 0.75 0.78 0.76 0.77 0.78 0.77
$\left[0.4,0.45\right]$ 0.95 0.99 0.98 0.58 0.95 0.87 0.71 0.78 0.75 0.75 0.79 0.77
$\left[0.5,1.0\right]$ 0.96 0.99 0.98 0.69 0.96 0.90 0.74 0.77 0.76 0.77 0.78 0.77
$\left[0.6,1.0\right]$ 0.97 0.99 0.99 0.61 0.96 0.97 0.75 0.77 0.76 0.78 0.78 0.76
$\left[0.7,1.0\right]$ 0.97 0.99 0.99 0.61 0.92 0.91 0.74 0.78 0.72 0.77 0.78 0.73
$\left[0.8,1.0\right]$ 0.97 0.99 0.99 0.63 0.92 0.96 0.74 0.78 0.73 0.76 0.78 0.73
$\left[0.9,1.0\right]$ 0.97 0.99 0.99 0.74 0.99 0.94 0.73 0.77 0.75 0.76 0.78 0.76
$\left[0.02,0.45\right]$ (ours) 0.98 0.0.99 0.0.99 0.93 0.96 0.95 0.82 0.83 0.83 0.84 0.84 0.85
Table 21: Impact of joint learning on closed-set classification and novelty detection on SUN397 across three novel groups. We compare three model variants: (i) Source-only, using only closed-set classification heads ($w^{c}$); (ii) ConOp, using only novelty detection heads ($w^{\alpha}$) learned via constrained optimization; and (iii) the full CoLOR model trained with a multitask objective for both classification and novelty detection. All methods are evaluated on $S_{{\mathcal{T}}}$ using ImageNet-pretrained ResNet50 features.
Metric Method Novel Classes (natural outdoor scenes/places)
$\alpha$ $0.06\pm 0.01$ $\alpha$ $0.06\pm 0.01$ $\alpha$ $0.08\pm 0.04$
AUROC ConOp $0.96\pm 0.02$ $0.97\pm 0.04$ $0.97\pm 0.02$ $0.96\pm 0.03$
AUPRC ConOp $0.77\pm 0.11$ $0.78\pm 0.21$ $0.87\pm 0.08$ $0.80\pm 0.14$
Top-1 Acc. Source-only $0.71\pm 0.03$ $0.73\pm 0.05$ $0.72\pm 0.03$ $0.72\pm 0.03$
CoLOR $\bm{0.77\pm 0.11}$ $\bm{0.78\pm 0.21}$ $\bm{0.87\pm 0.08}$ $\bm{0.80\pm 0.14}$
OSCR CoLOR $\bm{0.82\pm 0.03}$ $\bm{0.81\pm 0.03}$ $\bm{0.81\pm 0.05}$ $\bm{0.81\pm 0.04}$
Table 22: Impact of joint learning on closed-set classification and novelty detection on SUN397 across three novel groups. We compare three model variants: (i) Source-only, using only closed-set classification heads ($w^{c}$); (ii) ConOp, using only novelty detection heads ($w^{\alpha}$) learned via constrained optimization; and (iii) the full CoLOR model trained with a multitask objective for both classification and novelty detection. All the methods are evaluated on $S_{{\mathcal{T}}}$ and using pretrained CLIP ViT-L/14 features.
Metric Method Novel Classes (natural outdoor scenes/places)
$\alpha$ $0.06\pm 0.01$ $\alpha$ $0.06\pm 0.01$ $\alpha$ $0.08\pm 0.04$
AUROC ConOp $\bm{0.99\pm 0.01}$ $\bm{0.99\pm 0.00}$ $\bm{0.99\pm 0.00}$ $\bm{0.99\pm 0.01}$
AUPRC ConOp $\bm{0.97\pm 0.02}$ $\bm{0.96\pm 0.03}$ $\bm{0.97\pm 0.02}$ $\bm{0.97\pm 0.02}$
CoLOR $0.95\pm 0.04$ $0.95\pm 0.01$ $0.96\pm 0.03$ $0.95\pm 0.03$
Top-1 Acc. Source-only $0.97\pm 0.01$ $0.97\pm 0.01$ $0.96\pm 0.01$ $0.97\pm 0.01$
CoLOR $\bm{0.98\pm 0.01}$ $\bm{0.98\pm 0.00}$ $\bm{0.97\pm 0.01}$ $\bm{0.98\pm 0.01}$
OSCR CoLOR $\bm{0.96\pm 0.02}$ $\bm{0.97\pm 0.01}$ $\bm{0.96\pm 0.01}$ $\bm{0.96\pm 0.01}$
Figure 6: Synthetic experiment setup following the definition˜2 to measure novelty detection performance as a function of angular separation between $\mu$ and $\eta$ vectors, $\theta$, (Top) and $r_{\eta}$ (Bottom). We measure AUROC (Left) and AUPRC (Right). Here, $\alpha=0.15$, $d=3000$, $\sigma=1/d$, $r_{\mu}=0.1$. We train on 2000 samples and test on 3000 samples.
(a) AUROC vs $\theta$
(b) AUPRC vs $\theta$
(c) AUROC vs $r_{\eta}$
(d) AUPRC vs $r_{\eta}$
Table 23: Novelty detection performance as a function of angular separation between $\mu$ and $\eta$ vectors, $\theta$, in synthetic experiment setup following the definition˜2. Here, $\alpha=0.15$, $d=3000$, $\sigma=1/d$, $r_{\mu}=0.1$, $r_{\eta}=1.0$. We train on 2000 samples and test on 3000 samples.
$\theta$ $D$ CoLOR
AUROC AUPRC AUROC AUPRC
$\pi/2$ 0.51 0.12 0.94 0.65
$\pi/4$ 0.41 0.15 0.88 0.46
$\pi/8$ 0.27 0.22 0.69 0.21
$\pi/16$ 0.03 0.84 0.29 0.09
$\pi/32$ 0.00 0.99 0.06 0.07
Table 24: Novelty detection performance as a function of $r_{\eta}$ in synthetic experiment setup following the definition˜2. $\theta=\frac{\pi}{2}$ is the angular separation between $\mu$ and $\eta$ vectors. Here, $\alpha=0.15$, $d=3000$, $\sigma=1/d$, $r_{\mu}=0.1$. We train on 2000 samples and test on 3000 samples.
$r_{\eta}$ $D$ CoLOR
AUROC AUPRC AUROC AUPRC
1.0 0.51 0.12 0.94 0.65
0.9 0.50 0.12 0.93 0.62
0.8 0.49 0.13 0.92 0.58
0.7 0.46 0.13 0.91 0.53
0.6 0.43 0.14 0.89 0.47
0.5 0.44 0.14 0.86 0.39
0.4 0.34 0.17 0.81 0.31
0.3 0.21 0.27 0.70 0.21
0.2 0.11 0.46 0.46 0.12
0.1 0.01 0.96 0.09 0.07
A.7 Analysis of toy example
@@
A.8 Synthetic Experiment
To study the impact of separability between novel class and known classes, we conduct a synthetic experiment following the theoretical setup defined in definition 2 of the main draft as it allows us to explicitly control separability. In this study, we consider a simple Gaussian mixture setup and vary the angular separation between known and novel class means as well as the absolute distance of the novel class mean ($r_{\eta}$) to gradually vary the novel class separability. This enables us to directly examine how CoLOR’s performance changes as the separability assumption is increasingly violated. Please refer to fig.˜6, table˜23 and table˜24 for results and further details.